what would be your monthly car note if you bought a car for $24250, you paid off the loan in 5 years and interest on the loan was 4.25%?

use BA II Plus calculator!!
Answer #1

PV = -24250

N = 5*12 = 60

I/Y = 4.25%/12 = 0.3542 %

Find PMT

This is 449.34

Hence the monthly payment is $449.34